How will you compare it with the heat needed by alcohol with molar heat of vaporization 44.0 kJ at 35.0 g and boils at 82.5°C? Note that the molar mass of isopropyl alcohol is 60.1 g/mol.


1
Expert's answer
2022-03-04T15:21:02-0500

Let molar heat of vaporization be ∆Hvap

∆Hvap=40.7 KJ

Moles of H2O=50/18

=2.78 moles

Heat= 40.7×2.78

= 40.7×2.78

=113.146KJ